OpenStudy (anonymous):

Reduce fraction to lowest term. y+5/2y+10 1.) 1/4 2.) 1/2 3.) 2/5

OpenStudy (johnweldon1993):

*Hint..factor a 2 out of the denominator

OpenStudy (johnweldon1993):

\[\large \frac{y+ 5}{2(y + 5)}\] See anything you can simplify?

OpenStudy (anonymous):

Yes, the y+5

OpenStudy (johnweldon1993):

Right... \[\large \frac{\cancel{y + 5}}{2(\cancel{y + 5)}} = \frac{1}{2}\]

OpenStudy (anonymous):

Thank ya, sir..!:)

OpenStudy (johnweldon1993):

anytime :)